Find the most general antiderivative of the function f(x) -r-1 +5V2 22 4 cscx
Answer #1

- x 4 csc²x f(0) = + 5 a 22 - x + 4 cs.c²u sah 22 f(x) = e-3 -3/ t sa 4 CS C²u J fol dx = f ( x² + 5 232 - 4 csatu du - 3+1